DiabolikLover:

Nick can invest $10,000 in either one of two annuities. Annuity A has a 6% annual interest rate and requires a starting principal of $9,000, plus annual $100 deposits for the next 10 years. Annuity B has a 12% annual interest rate and requires a starting principal of $9,000, plus annual $200 deposits for the next 5 years.

DiabolikLover:

Mage:

Let's start by cutting it into parts. What we're doing with this equation is finding both values of annuities A and B and subtracting them. Let's begin with annuity A. First, take your numbers out of the word problem: $10,000 for investment 6% annual interest rate $9,000 starting principal $100 deposits annually for 10 years

DiabolikLover:

ok got that, whats thr right formula to find out the final. i thought i found one but idk where the investment part goes in the equation

Mage:

PV = P x (1 - (1 + r ) ^- n) / r PV = present value of an ordinary annuity P = value of each payment r = interest rate per period n = number of periods

DiabolikLover:

ok i tried that one before and got 91.333... which isnt possible

Mage:

how did you substitute into the equation?

DiabolikLover:

100 x (1 - (1+6)^-10) /6 = 99.999999941

DiabolikLover:

it just got even more confusing idk where the 9,000 or the 10,000 goes

DiabolikLover:

Cheese it, im guessing this is way too confuisng and im just waisting your time at this point

DiabolikLover:

thx for the help tho

Mage:

nono, hold on

Mage:

PV would be the 9,000 because it is the initial deposit

DiabolikLover:

ok still not understanding it unforuetly

Mage:

i just realized

DiabolikLover:

what

Mage:

i feel so dumb. $9000 is the principal sort of like the down payment, so we go ahead and take it off of the $10000, leaving us with $1000 with 10 years of $100 annual payments and a 6%, or .06 rate of return and now we can calculate

Mage:

PV = P x (1 - (1 + r ) ^- n) / r, right? so if PV = present value of an ordinary annuity P = value of each payment r = interest rate per period n = number of periods $1000 = 100 x (1 - (1 + 6%)^-10) / 6%

DiabolikLover:

yee that makes more sense

Mage:

i'm sorry! i feel so slow rn

DiabolikLover:

lol its fine, i guessed and still made it out with a good grade. if its night there you should go sleep lol

Mage:

nO i wAnT tO sOLvE!! sleep is for casuals